Obtain the inverse demand and supply functions if demand and supply of Chinese coal are
Qd = 60 − 0.5P.
Qs = −2 + 2P.
1
Expert's answer
2020-11-10T10:35:11-0500

"P=\frac{60-Qd}{0.5}=120-2Qd"

"P=0.5Qs+1"